# Semantic-guided Search for Efficient Program Repair with Large Language Models

### v1

@@ 2.1.2. Studied LLM-based Program Repair techniques
In this study, we also investigated the potential of memory reduction techniques through engineering efforts and their impact on our findings. These techniques include:
  • Quantization: A widely used technique to reduce the memory usage of LLMs by converting their weights from high-precision (32-bit) to lower-precision values, such as 8-bit or 4-bit. In this work, we apply 4-bit quantization, reducing the LLM’s weights from 32-bit to 4-bit.
  • Sequential Beam Search: An implementation variant of beam search proposed by Saibo Geng (developers, 2024). This approach processes each beam sequentially through the model before concatenating them into a single output.
@@
We implement LLM-based APR techniques using Python, leveraging the PyTorch and HuggingFace libraries for efficient execution. To enhance the accuracy of our implementations and minimize potential bias, we collected and integrated the inference code, including repair prompts and trained models, from their original repositories into a unified framework. All implementations utilize HuggingFace’s standard beam search, a widely used method in LLM-based techniques. We perform inference with a batch size of one, i.e., trying to repair one buggy program at a time. All experiments are conducted on a single NVIDIA H100 GPU with 80GB of VRAM, a state-of-the-art and widely-used hardware for Deep Learning and Large Language Models.
@@ 2.2. RQ1: Impact of Beam Size on Memory Efficiency
Figure 1. Memory usage of LLM-Based APR techniques across various beam sizes on an NVIDIA A-100 GPU with 80 GB VRAM, using both original and quantized Models. ‘O‘ and ‘Q‘ denote original and quantized models, respectively. ‘BS‘ and ‘SeqBS‘ denote beam search and sequential beam search, respectively. RepairLlama with quantization and sequential beam search were not assessed because (1) the original models were already quantized and (2) sequential beam search does not support this architecture.
Figure 1 illustrates the experimental results of RQ1. Overall, there is a significant increase in VRAM usage as the beam size expands from 0 to 200. For instance, in CodeGen-2B, average memory usage jumps from approximately 20GB at a beam size of 10 to around 70GB at a beam size of 200. It is important to note that our figures are capped at 80GB due to hardware limitations, representing only the lower bound of the actual numbers, which may be much higher. Similarly, peak memory usage climbs dramatically from about 30GB to 80GB, which is the maximum our hardware can support. These trends are consistent across other LLM-based APR techniques. Notably, this increase in VRAM usage induces a substantial rise in crashes due to OOM errors. For example, the OOM ratio in CodeGen-2B increases from nearly negligible at smaller beam sizes to about 60% at a beam size of 200. Larger models, such as CodeGen-6B and InCoder-6B, show even higher OOM crash rates, exceeding 80
Next, we explore the potential of 4-bit quantization and sequential beam search, which are widely used to optimize model memory requirements, to reduce VRAM usage. Figure 1 shows that quantization significantly reduces both average and peak VRAM usage, decreasing the frequency of OOM crashes. However, sequential beam search does not notably reduce the memory usage of LLM-based APR techniques and even increases the frequency of OOM crashes. Sequential beam search involves trade-offs and does not guarantee reduced memory usage in all scenarios. This observation has been also confirmed by the authors of sequential beam search. Further details of this can be found in our online Appendix (Authors, 2023).
More importantly, we also observe that VRAM usage for LLM-based APR techniques with both quantization and sequential beam search still increases substantially as beam size increases. For instance, average VRAM usage for CodeGen-2B rises from less than 10GB at a beam size of 10 to around 60GB at a beam size of 200. Its peak memory usage also quickly approaches the hardware limit of 80GB. Consequently, even with memory reduction techniques, LLM-based APR techniques continue to experience a significant number of OOM crashes in our benchmark dataset, ranging from 30% to 75% at a beam size of 200.
Answers to RQ1: Both average and peak memory usage of LLM-based APR techniques rise significantly as the beam size increases, inducing up to 80% out-of-memory crashes, on a NVIDIA A100 with 80GB VRAM. Even with memory reduction techniques, memory usage still escalates and crash rates remain substantial as the beam size increase.
@@ 2.3. RQ2: Impact of Beam Size on Effectiveness
Table 1. Effectiveness of LLM-based APR techniques was evaluated across different beam sizes on an NVIDIA A100 GPU with 80 GB VRAM, using both original and quantized models. ‘O‘ and ‘Q‘ denote original and quantized models, respectively. ‘BS‘ and ‘SeqBS‘ denote beam search and sequential beam search, respectively. RepairLlama with quantization and sequential beam search were not assessed because (1) the original models were already quantized and (2) sequential beam search does not support this architecture.
Model O + BS Q + BS O + SeqBS
10 25 50 100 200 10 25 50 100 200 10 25 50 100 200
CodeGen-2B 66 80 87 75 38 64 80 98 107 89 65 73 80 70 27
CodeGen-6B 72 93 75 39 11 65 95 114 91 55 62 91 74 39 5
InCoder-1B 74 108 119 120 96 71 102 124 134 126 69 104 107 109 65
InCoder-6B 77 95 92 61 27 77 100 123 114 83 75 88 76 34 14
RepairLlama 139 169 160 134 75 _ _ _ _ _ _ _ _ _ _
Table 1 present our experimental results for RQ2. We can see that the number of plausible patches, only shows an upward trend at smaller beam sizes and remarkably declines at larger sizes. For instance, the number of plausible patches generated by RepairLlama increases from 139 at a beam size of 10 to 169 at 25, then drops significantly to 134 at 100 and further to 75 at 200, although the ratio of plausible patches remains at 0.60. This reduction is primarily due to the high frequency of OOM crashes, which range from 30-60% at larger beam sizes, as depicted in Figure 1. These results highlight the significant impact of LLMs’ extensive memory usage on their performance with large beam sizes.
Next, similar to RQ1, we examine the impact of 4-bit quantization and sequential beam search on memory usage and the effectiveness of LLM-based APR techniques. Overall, the effectiveness of LLM-based APR techniques remains consistent with sequential beam search, maintaining the same trends. Meanwhile, quantization sustains an upward trend in effectiveness up to a beam size of 100. For example, the number of plausible patches generated by the quantized CodeGen-2B increased from 64 at a beam size of 10 to 107 at a beam size of 100, surpassing the original model by 42.7% at the same beam size. This improvement is due to a reduction in OOM crashes, from 30% to less than 5%. However, despite these gains, the effectiveness of quantized models still declines at larger beam sizes, such as 200, similar to the trend seen in the original models. For instance, the effectiveness of the quantized CodeGen-2B dropped from 114 plausible patches at a beam size of 50 to 91 at 100, and further to 55 at 200, reflecting declines of 20% and 48%, respectively.
Answers to RQ2: As beam size increases, the effectiveness of LLM-based APR generally rises, peaking at mid-range beam sizes, e.g., 25 or 50, before declining at larger sizes due to higher OOM crash rates, particularly in larger models. Even with memory reduction techniques, the APR’s effectiveness still quickly degrades at beam sizes larger than 100.
@@ 3. Approach
Our findings in Section 2 highlight that scaling beam size to a certain threshold may help improve the effectiveness of APR at the cost of extensive memory consumption, above which the effectiveness quickly degrades even on state-of-the-art hardware.
To address these limitations, we propose a novel approach, FLAMES, that combines LLM-based and search-based APR for an efficient, semantic-guided patch generation. At high level, FLAMES initially generates candidate patches using greedy decoding (i.e., beam search with a beam size of 1) and then leverages semantic feedback from test validation and a specialized best-first search, namely PG-TD (Zhang et al., 2023a), to iteratively explore potential candidate patches. FLAMES uses a small beam size to keep memory usage low, while allowing it to explore a greater number of potential candidate via the guidance of semantic feedback with best-first search.
Figure 2 illustrates the overall pipeline of FLAMES. The search algorithm of FLAMES guides the patch generation process in LLM-based APR through four main phases: Selection, Expansion, Simulation, and Backpropagation. In this section, we first re-formulate the patch generation task in LLM-based APR as a search problem. Then, we present details of these four stages used to generate candidate patches for a given buggy program.
@@
3.1. Problem Formulation
We first formulate the patch generation process in APR as a Markov Decision Process (Sutton and Barto, 2018) which contains state, action, transition function and reward function as inspried by prior works (Ellis et al., 2019; Zhang et al., 2023a).
State. The state is a triplet $s=(pp,spec,bp)$, where $bp$ is the buggy program which needs to be fixed, $pp$ is a partial patch for $bp$ represented by a sequence of tokens and $spec$ is a specification which describes desired behaviors of $bp$. The initial state is $(\langle sos\rangle,spec,bp)$ where $\langle sos\rangle$ denotes the start of sequence token in LLMs.
Action. An action $a$ is a token from the vocabulary of fine-tuned LLM for APR. We also have special tokens, i.e., terminal tokens, that indicates that a program is complete. These terminal tokens are pre-defined for each LLM. We define that a partial patch $pp$ is complete if and only if the last token of $pp$ is a terminal token.
Transition function. The transition $T$, takes the set of partially-contructed patch $pp$ and applies the action $a$ as follows:
(1) $T(pp,a)=pp\oplus[a]$
where $\oplus$ denote a sequence concatenation.
Reward. The reward of a state $s=(pp,spec,bp)$ is defined to be either:
  • 0 if $pp$ is a partial patch;
  • $R(pp,spec)$ if $pp$ a complete patch, where $R$ is a function which measure how likely $pp$ statisfies $spec$.
Intuitively, $R$ serves for same purpose as objective function in search-based APR (Le Goues et al., 2011; Le et al., 2016).
3.2. Overview
To efficiently and effectively search for repairs, FLAMES maintains a tree structure where each node represents a token and the path from the root to each node represents a partially constructed patch. Then, FLAMES starts from an empty patch and search for a complete patch with optimal reward. The process includes four stages: (1) selection phase for selecting the most promising partial patch, i.e., branch of tree, to visit using a policy algorithm (Section 3.3); (2) expansion phase for expanding the search tree by adding potentially tokens (Section 3.4); (3) evaluation phase for estimating the reward of current state using simulation (Section 3.5); and (4) backpropagation phase for updating value of partial patches that their potential (Section 3.6). We elaborate details of the stages below.
3.3. Selection
At this stage, our goal is to select the most promising partial patch to explore. To achieve this, we employ a policy algorithm that strikes a balance between exploiting potentially high-reward states and exploring less-visited ones. Specifically, an action (i.e., adding a token) is chosen based on three criteria: (1) the action’s historical reward in relation to the current state is high, indicating strong potential; (2) language models (LLMs) predict that the action is a suitable next token, as reflected by its prediction probability; and (3) the resulting state after taking the action is relatively under-explored.
To implement this approach, we evaluate three commonly-used policy algorithms in Monte Carlo Tree Search (MCTS): UCB (Kocsis and Szepesvári, 2006), and two variants of P-UCB (Silver et al., 2017)—one with fixed weights (Fixed P-UCB) and the other with variable weights (Variable P-UCB). Our results indicate that Variable P-UCB is the most effective algorithm for the APR task, leading us to adopt it as our primary policy algorithm (See Section 4.6 for further details.)
@@ 3.4. Expansion
This stage aims to expand the search tree by adding possible next tokens. As the number of possible tokens in vocabulary is huge, a random sampling, which is used in conventional MCTS (Kocsis and Szepesvári, 2006) may likely return in invalid and low-potential tokens. Therefore, FLAMES leverages fine-tuned LLM to suggest top-$k$ next tokens with the highest prediction score given the current state following PG-TD. Note that, as a state can be re-visited multiple times during the tree search, FLAMES also leverages a caching mechanism, namely tree structure caching, to avoid redudant computation. In particular, whenever LLMs make a prediction, the input sequence, i.e., state and top-$k$ next tokens is cached using a hash map. Then, when the algorithm needs LLM to predict top-$k$ next tokens, such information can be reused if it exists in the cache. In this work, we set $k=10$ and further discuss the impact of $k$ on our method in Sections 4.6.2.
@@ 3.5. Evaluation
This stage aims to evaluate the potential of the current state, i.e., how likely the partial patch leads to a optimal (complete) patch. Similar to expansion, it is impossible to evaluate the current state using an random simulation as used in conventional MCTS algorithm. Consequently, FLAMES leverages beam search to automatically complete the current state, i.e., a partial patch $pp$, to form a complete patch $cp$. However, it is noteworthy that leveraging the beam search as in original PG-TD still hinders memory efficiency. Therefore, instead of beam search, FLAMES utilizes a greedy search, that is equivalent to a beam search with a beam of size 1, to find the complete patch for current state while keeping low memory consumption.
Then, given the completed patch, FLAMES estimates the potential of current state by calculating a reward function $R(cp,spec)$. Theoretically, $spec$ and reward function $R$ can takes any kinds of specification and a corresponding pre-defined reward function. In this work, following common practices in literature (Gao et al., 2022), we leverage developer-written test cases as our specification and utilize the passing ratio, which is widely-used in search based APR (Gao et al., 2022), as our reward function. More formally, given a program $cp$ and a set of test cases $spec$, we define our reward function as follows: $R(cp,spec)=\dfrac{f_{pass}}{f_{pass}+f_{fail}}$ with $f_{pass}$ and $f_{fail}$ is the number of passing and failing test cases observed when running $cp$ on $spec$.
@@ 3.6. Backpropagation
This stage aims to update the value of nodes in the tree for guiding next steps of the search process. Intuitively, the stage allows FLAMES to provide feedback from test validation for guiding the patch generation process of LLMs. Particularly, the reward of a state (calculated based on automatically-completed program) is backpropagated to its parents recursively until the root is reached. For all state and action pair $(s^{\prime\prime},a^{\prime\prime})$ along the path from current state to root node, FLAMES updates their Q-value as follows:
(2) $Q(s^{\prime\prime})=max(Q(s^{\prime\prime}),R(cp,spec))$
where $R(cp,spec)$ is reward value of current state - action pair $(s,a)$ with $cp$ is automatically-completed patch of $s$.
@@
We implemented FLAMES by extending the PG-TD (Zhang et al., 2023a) framework, based on the DynaGym toolkit. It is developed in Python using PyTorch, HuggingFace, and OpenAI Gym. We use RepairLLama, the best-performing model among LLM-based repair techniques, as our default LLM-based repair model. For repair, FLAMES generates up to 200 patches and sets a 30-minute timeout for each bug, as recommended by developers (Noller et al., 2022). Experiments are conducted on an NVIDIA A100 GPU with 80GB VRAM, 250 GB RAM, and a 32-core Intel Xeon CPU at 2.90GHz, running Red Hat Enterprise Linux 9.3 and Java 1.8.0_241. This setup is comparable to those used in related works (Silva et al., 2023; Wei et al., 2023).

### v2

@@ 2.1.2. Studied LLM-based APR techniques
In this study, we also investigate the potential of memory reduction techniques, including quantization and sequential beam search, through engineering efforts and their impact on our findings. Quantization is a widely used technique to reduce the memory usage of LLMs by converting their weights from high- to lower-precision values. In this work, we applied 4-bit quantization, reducing the LLM’s weights from 32 bits to 4 bits. For clarity, we denote the original LLM-based APR as full-precision, and its quantized counterpart as quantized in the following sections. Meanwhile, sequential beam search is a variant of the implementation of the beam search proposed by Saibo Geng (developers, 2024), which processes each beam sequentially through the model before concatenating them into a single output.
@@
We implemented LLM-based APR techniques in Python using PyTorch and HuggingFace. To improve reliability and reduce bias, we integrated inference code, repair prompts, and trained models from the original repositories into a unified framework. Our implementations also used HuggingFace’s standard beam search. Inference was performed with a batch size of one, processing one buggy program at a time. All experiments were run on a single NVIDIA A100 GPU (80GB VRAM).
@@ 2.2. RQ1: Impact of Beam Size on Memory Efficiency
Figure 1(a) illustrates the experimental results of RQ1. Overall, we can see that VRAM usage increases significantly as the beam size increases from 0 to 200 in the full-precision versions of five studied techniques. For instance, in InCoder-1B, the average memory usage substantially increases from 20GB at a beam size of 10 to around 60GB at a beam size of 200. Similarly, peak memory usage increases dramatically, reaching maximum memory limits at a beam size of only 50. Notably, this increase in VRAM usage leads to a substantial increase in crashes due to Out-of-Memory (OOM) errors. For example, the OOM ratio in InCoder-1B increases from nearly negligible at smaller beam sizes to approximately 50% at a beam size of 200. Larger models, e.g. CodeGen-6B and InCoder-6B, show even higher OOM rates, nearly 80%, respectively.
(a) Memory Usage
(b) Effectiveness
Figure 1. Memory usage and effectiveness of LLM-based APR techniques across beam sizes on an NVIDIA A100 (80GB) using full-precision (FP) and quantized (Q) models. BS and SeqBS denote standard and sequential beam search.
Next, we investigated 4-bit quantization and sequential beam search, commonly used to optimize model memory requirements, for reducing VRAM usage. As shown in Figure 1(a), quantization significantly reduces the average and peak VRAM consumption, thereby reducing OOM crashes. However, sequential beam search presents trade-offs and does not consistently decrease memory usage, as confirmed by its authors and our theoretical analysis (for details, please see our online Appendix (Authors, 2023)). More importantly, even with these techniques, VRAM usage for LLM-based APR techniques continues to increase with increasing beam size. For example, CodeGen-2B’s average VRAM usage rises from less than 10GB at a beam size of 10 to 60GB at 200, with peak usage nearing the 80GB hardware limit. As a result, despite memory reduction techniques, LLM-based APR techniques still face frequent OOM crashes, occurring in 30–85% of cases with a beam size of 200.
Answers to RQ1: Both the average and peak memory usage of LLM-based APR techniques increase significantly as the beam size increases, inducing up to 80% out-of-memory crashes on an A100 with 80GB of VRAM. Even with memory reduction techniques, memory usage still escalates, and crash rates remain substantial as the beam size increases.
@@ 2.3. RQ2: Impact of Beam Size on Effectiveness
Figure 1(b) presents our experimental results for RQ2. Overall, we can see that the number of plausible patches generated by full-precision versions of LLM-based APR only shows an upward trend at smaller beam sizes and remarkably declines at larger sizes. For instance, the number of plausible patches generated by RepairLlama increases from 139 at a beam size of 10 to 169 at 25, then drops significantly to 134 at 100 and further to 75 at 200, although the ratio of plausible patches remains at 0.60. This reduction is primarily due to the high frequency of OOM crashes, which range from 30-60% at larger beam sizes, as depicted in Figure 1(a). These results highlight the significant impact of LLMs’ extensive memory usage on their performance with large beam sizes.
Next, similar to RQ1, we examined the impact of 4-bit quantization and sequential beam search on memory usage and the effectiveness of LLM-based APR techniques. Overall, the effectiveness of LLM-based APR techniques remains consistent with sequential beam search, maintaining the same trends. Meanwhile, quantization maintains an upward trend in effectiveness up to a beam size of 100. For example, the number of plausible patches generated by quantized CodeGen-2B increased from 64 at a beam size of 10 to 107 at a beam size of 100, surpassing the original full-precision model by 42.7% at the same beam size. This improvement is due to a reduction in OOM crashes, from 30% to less than 5%. However, despite these gains, the effectiveness of quantized models still declines at larger beam sizes, such as 200, similar to the trend seen in the full-precision models. For example, the effectiveness of the quantized CodeGen-2B dropped from 114 plausible patches at a beam size of 50 to 91 at 100, and further to 55 at 200, reflecting declines of 20% and 48%, respectively.
Answers to RQ2: As the beam size increases, the effectiveness of LLM-based APR increases, peaking at mid-range beam sizes before declining at larger sizes due to higher OOM crash rates, particularly in larger models. Even with memory reduction techniques, the APR’s effectiveness still quickly degrades at beam sizes larger than 100.
@@ 3. Approach
Our findings in Section 2 highlight that scaling the beam size to a certain threshold may help improve the effectiveness of APR at the cost of extensive memory consumption, above which the effectiveness quickly degrades even on state-of-the-art hardware. To address these limitations, we propose a novel approach, FLAMES, which combines LLM-based and search-based APR for memory-efficient, semantic-guided patch generation.
Figure 2 illustrates the overall pipeline of FLAMES, which employs PG-TD as the core algorithm to guide patch generation in LLM-based APR. PG-TD is a best-first search algorithm, inspired by Monte Carlo Tree Search (Kocsis and Szepesvári, 2006), to decode Transformer models. PG-TD formulates the decoding process of Transformers models as a search problem on a tree structure whose nodes represent tokens. PG-TD then iteratively expands this tree with the information provided by the models and searches for an optimal candidate using a predefined reward function. In the following sections, we first present a high-level overview of patch generation of LLM-based APR with PG-TD and then present a detailed methodology of PG-TD’s steps in APR context.
3.1. Patch Generation with PG-TD
Algorithm 1 outlines the patch generation process in APR using the PG-TD algorithm. At a high level, PG-TD conducts an iterative search for plausible patches. This search proceeds by refining partial repair candidates over multiple iterations. At each iteration, PG-TD selects the most promising partial patch from the current search space, as described in Section 3.2. The chosen partial patch is then used in two key operations. First, since it is the most promising candidate, PG-TD will further expand the search space by incorporating most of the next potential tokens for the partial patch, predicted by an LLM-based APR model (Section 3.3). Second, the partial patch is simulated into a complete patch candidate using the same LLM-based APR model and is subsequently evaluated against a predefined set of correctness specifications to compute a reward value (Section 3.4). This reward is then backpropagated through the search tree to update the estimated potential of the nodes in the partial patch (Section 3.5). This iterative process continues until a plausible patch is discovered, that is, a candidate that satisfies the given specifications or the iteration limit is reached.
3.2. Selection
At this stage, our goal is to select the most promising partial patch to explore. To achieve this, we employ a policy algorithm that strikes a balance between exploiting potentially high-reward states and exploring less-visited ones. Specifically, an action (i.e., adding a token) is chosen based on three criteria: (1) the action’s historical reward in relation to the current state is high, indicating strong potential; (2) language models (LLMs) predict that the action is a suitable next token, as reflected by its prediction probability; and (3) the resulting state after taking the action is under-explored.
To implement this approach, we evaluate three commonly used policy algorithms in Monte Carlo Tree Search (MCTS): UCB (Kocsis and Szepesvári, 2006), and two variants of P-UCB (Silver et al., 2017), one with fixed weights (Fixed P-UCB) and the other with variable weights (Variable P-UCB). Our results indicate that Variable P-UCB is the most effective algorithm for the APR task, leading us to adopt it as our primary policy algorithm (See Section 4.6 for further details.)
@@ 3.3. Expansion
This stage aims to expand the search tree by adding possible next tokens. As the number of possible tokens in the vocabulary is huge, random sampling, which is used in MCTS (Kocsis and Szepesvári, 2006) is likely to return in invalid and low-potential tokens. Therefore, FLAMES leverages fine-tuned LLM to suggest top-$k$ next tokens with the highest prediction score given the current state following PG-TD (Zhang et al., 2023a). Note that, as a state can be re-visited multiple times during the tree search, FLAMES also leverages a caching mechanism to avoid redundant computation. In particular, whenever LLMs make a prediction, the input sequence, i.e., state and top-$k$ next tokens, is cached using a hash map. Then, when the algorithm needs LLM to predict top-$k$ next tokens, such information can be reused if it exists in the cache. In this work, we set $k=10$ and further discuss the impact of $k$ on our method in Sections 4.6.2.
Algorithm 1 Semantic-guided Patch Generation with PG-TD. SELECTSELECT, EXPANDEXPAND and BACKPROPBACKPROP are presented in Section 3.2,  3.3 and  3.5, respectively. SIMULATESIMULATE and EVALEVAL are presented in Section 3.4.
0:bpbp: buggy program; $\mathcal{L}$: LLM-based Program Repair model; $\mathcal{S}$: set of correctness specifications; max_itersmax\_iters: number of iterative refinement steps
0:cpcp: plausible patch for bpbp or $\mathit{None}$ if no patch found
[…]
3:for $i=1$ to max_itersmax\_iters do
4:  $pp\leftarrow SELECT(root,reward)$ $\triangleright$ Select best partial patch
5:  $root\leftarrow EXPAND(root,\mathcal{L},bp)$ $\triangleright$ Expand search tree
6:  $cp\leftarrow SIMULATE(pp,\mathcal{L},bp)$ $\triangleright$ Simulate complete patch cpcp
7:  $r\leftarrow EVAL(cp,\mathcal{S},bp)$ $\triangleright$ Evaluate rewards of pppp using cpcp
8:  if $r\neq 1$ then
9:   $reward\leftarrow BACKPROP(reward,pp,r)$ $\triangleright$ Update rewards
10:  else
11:   return cpcp $\triangleright$ Return plausible complete patch
12:  end if
13:end for
14:return $\mathit{None}$
@@ 3.4. Evaluation
This stage aims to evaluate the potential of the current state, i.e., how likely the partial patch leads to a optimal (complete) patch. Similar to expansion, it is impossible to evaluate the current state using an random simulation as used in MCTS algorithm. Consequently, FLAMES leverages beam search to automatically complete the current state, i.e., a partial patch $pp$, to form a complete patch $cp$. However, it is noteworthy that leveraging the beam search as in original PG-TD still hinders memory efficiency. Therefore, instead of beam search, FLAMES utilizes a greedy search that is equivalent to a beam search with a beam of size 1, to find the complete patch for the current state while keeping low memory consumption.
Then, given the completed patch, FLAMES estimates the potential of current state by calculating a reward function $R(cp,spec)$. Theoretically, specspec and reward function $R$ can takes any kinds of specification and a corresponding pre-defined reward function. In this work, following common practices in literature (Gao et al., 2022), we leverage developer-written test cases as our specification and utilize the passing ratio, which is widely-used in search based APR (Gao et al., 2022), as our reward function. More formally, given a program $cp$ and a set of test cases specspec, we define our reward function as follows: $R(cp,spec)=\frac{f_{pass}}{f_{pass}+f_{fail}}$ with $f_{pass}$ and $f_{fail}$ is the number of passing and failing test cases observed when running $cp$ on specspec.
@@ 3.5. Backpropagation
This stage aims to update the value of nodes in the tree for guiding next steps of the search process. Intuitively, the stage allows FLAMES to provide feedback from test validation for guiding the patch generation process of LLMs. Particularly, the reward of a state (calculated based on automatically-completed program) is backpropagated to its parents recursively until the root is reached. Similar to other reinforcement learning algorithms (Sutton and Barto, 2018; Wiering and Van Otterlo, 2012; Kaelbling et al., 1996), we aim to maintain the maximum observed reward as an expected return value for a given state $s$ and the corresponding action $a$.
To do so, for all state and action pair $(s,a)$ along the path from the current state to the root node, FLAMES updates their Q-value by $Q(s,a)=max(Q(s,a),R(cp,spec))$, where $R(cp,spec)$ is the reward value of the current state action pair $(s,a)$ with $cp$ is an automatically complete patch of $s$.
@@
We implemented FLAMES by extending the PG-TD (Zhang et al., 2023a) framework based on the DynaGym toolkit. We use RepairLLama, the best-performing model among LLM-based repair techniques, as our default LLM-based repair model. For repair, FLAMES generates up to 200 patches and sets a 30-minute timeout for each bug, as recommended by developers (Noller et al., 2022). Experiments are conducted on an NVIDIA A100 GPU with 80 GB VRAM, 250 GB RAM, and a 32-core Intel Xeon CPU at 2.90 GHz, running Red Hat Enterprise Linux 9.3 and Java 1.8.0_241. This setup is comparable to those used in related works (Silva et al., 2023; Wei et al., 2023).